Find the critical value to test the claim that μ1 < μ2. Two samples are random, independent, and come from populations that are normal. The sample statistics are given below. Assume that σ 2/1= σ2/2. Use α = 0.05.

n1 = 15 n2 = 15 x1 = 25.74 x2 = 28.29 s1 = 2.9 s2 = 2.8
Answer #1

Solution :

Given that ,

Degrees of freedom = n1 + n2 - 2 = 15 + 15 - 2 = 28

t 0.05,28 = -1.701

Critical value = -1.701